find volume of solid generated by revolving the region bounded by y = √(x) and the lines y = 3 and x = 0 about the x-axis
1
Expert's answer
2020-12-17T19:12:42-0500

The formula of finding the volume of the solid generated by revolving the region about x axis is found by multiplying π by the subtraction of integrations of the square of the functions in given interval ;

V="\int" π *(f1(x)2-f2(x)2)dx. after sketching the graph of the function we can find that the boundary of the solid in x axis [0;9].

integrating in this interval,

V= "\int"09 π*((32)-x)dx= π"\int"0 9(9-x)dx= π*(9x-x2/2)]09π*81/2;

the volume of the solid is V=π *81/2